Five Measures of Solvency or Profitability

The balance sheet for Garcon Inc. at the end of the current fiscal year indicated the following:

Bonds payable, 6% $1,900,000
Preferred $10 stock, $50 par 73,500
Common stock, $12 par 308,700.00

Income before income tax was $319,200, and income taxes were $47,250 for the current year. Cash dividends paid on common stock during the current year totaled $54,023. The common stock was selling for $70 per share at the end of the year.

Determine each of the following. Round answers to one decimal place, except for dollar amounts which should be rounded to the nearest whole cent. Use the rounded answers for subsequent requirements, if required.

a. Times interest earned ratio times
b. Earnings per share on common stock $
c. Price-earnings ratio
d. Dividends per share of common stock $
e. Dividend yield %
Answer #1

Interest expense = 1900000*6% = 114000

a) Times interest earned ratio = earning before interest and tax/Interest expense = (319200+114000)/114000 = 3.8

b) Earning per share on Common Stock = (Net income-Preferred dividend)/Share outstanding = (319200-47250-14700)/25725 = 10 per share

c) Price earning ratio = MPS/EPS = 70/10 = 7

d) Divided per share = 54023/25725 = 2.10 per share

e) Dividend yield = DPS/MPS = 2.10/70 = 3%